OpenStudy (anonymous):

Hey , can you guys help me with vectors. consider the displacment vectors A=i+6j meter , B=3!-7j meter and C=6i-2j meter 1) if aA-bB =2C what are the values of a and b 2) if A-B-4D=0 , what are the components of D?

OpenStudy (goformit100):

A warm Welcome to OpenStudy. I can help guide you through this useful site. You can ask your questions to me or you can message me. Please use the chat for off topic questions. Remember to give the person who helped you a medal by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it.

OpenStudy (ikram002p):

its easy just start like this part a aA-bB=2C a(i+6j)-b(3i+7j)=2(6i+2j) ai-3bi=12i 6aj-7bj=4j now u have two equations a-3b=12 6a-7b=4 \[b=\frac{ -32 }{ 11 }\] find a

OpenStudy (anonymous):

Thank you so much man , I posted this question early because I thought I won't get an answer that fast , I'll check what you gave me when I start studying, And i'll update when There's a point that isn't clear. Thanks a lot ^.^